# Alignment-Sensitive Minimax Rates for Spectral Algorithms with Learned Kernels

### v2

@@ 3.2 Span Profile
The span profile $\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}}$ is a well-defined object that depends only on $\bm{\theta}^{*}$ and the ordering of $\bm{\lambda}$, and it summarizes how $\sigma^{2}$ affects the ESD. Theorem˜3.2 suggests that for two spectra $\bm{\lambda}^{(1)}$ and $\bm{\lambda}^{(2)}$, we can compare their alignments with the signal by the ratio of $r(\tau)={\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}^{(1)}}(\tau)}/{\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}^{(2)}}(\tau)}$ for small $\tau$, because, if this ratio is very small (and in particular if the limit is 0 for $\tau\to 0$), then a kernel method using $\bm{\lambda}^{(1)}$ can achieve a smaller risk than one that uses $\bm{\lambda}^{(2)}$. Such comparisons are not as convenient in classical theory. See Section˜9 for more illustrations.
A closely related object is the trade-off function of $\bm{\theta}^{*}$ relative to $\bm{\lambda}$, which is defined as
$\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}}(k)\;=\;\frac{1}{k}\sum_{i=k+1}^{d}\left(\theta_{\pi_{i}}^{*}\right)^{2}\;=\;\frac{1}{k}\sum_{\,i:\,\lambda_{i}<\lambda_{\pi_{k}}\,}\left(\theta_{i}^{*}\right)^{2},k\in[d].$ (10)
@@
Property (3) in Proposition˜3.5 suggests that the faster $\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}}(\cdot)$ decreases, the better the spectrum $\bm{\lambda}$ aligns with the signal $\bm{\theta}^{*}$. In the extreme case where the ordering of $\lambda_{i}$ matches the ordering of $|\theta_{i}^{*}|^{2}$}, the decay of $\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}}(\cdot)$ is the fastest, which leads to the most favorable span profile.
@@ 4 Minimax Optimal Convergence Rates
(1) Suppose $K_{n}\asymp n^{a}$ where $0<a<1$. For any $k$, we have $L_{k}\asymp M_{k}\asymp k^{1/a}$. Since $k/k^{1/a}$ is decreasing, ˜4.1 holds.
(2) Suppose $K_{n}\asymp(\log n)^{b}$ where $b>0$. For any $k$, we have $L_{k}\asymp M_{k}\asymp e^{k^{1/b}}$. Since $k/e^{k^{1/b}}$ is decreasing, ˜4.1 holds.
The next theorem provides a lower bound on the minimax risk in Equation˜12.
@@ 5 Adaptive Eigenvalues via Over-parameterized Gradient Flow
This section will investigate the benefits of learning eigenvalues via over-parameterized gradient flow (OP-GF) in sequence models [28] through the lens of ESDs. Section˜9.2 outlines the framework to explain the benefits of adaptive kernels and presents an experiment using a linear network. Analyzing eigenfunction evolution theoretically is more difficult and is left for future work.
@@
Suppose that ˜5.1 holds and the initialization in Equation˜13 is $b_{0}=c_{B}D^{\frac{D+1}{D+2}}\varepsilon^{\frac{1}{D+2}}$. Define $t_{2}=C\cdot D^{\frac{D}{D+2}}(\varepsilon)^{-\frac{2D+2}{D+2}}$. There exist some constants $c$, $C$, $C_{M}$, $C_{max}$, $C_{\eta}$, $c_{\eta}$, $c_{B}$, and $c^{\prime}$, such that with probability larger than 14/n1-4/n, we have
@@
Theorem˜5.2 shows that OP-GF reduces the ESD. The conditions ensure that strong signals with small initial eigenvalues have larger adapted eigenvalues than weak signals.
@@ 6 Numerical experiments
Data Generation. We utilize the misalignment setting in Li and Lin [28] to specify a $d$-dimensional sequence model. We fixed the eigen-decay rate $\gamma>0$, the signal decay rate $p>0$, and the number of nonzero signals $J$. Given any misalignment parameter $q\geq 1$, we set eigenvalues as $\lambda_{j}=j^{-\gamma},j\in[d]$, and set the true nonzero parameters as $\theta_{\ell(j)}^{*}=C\cdot j^{-\frac{p+1}{2}}$, where $\ell(j)=[j^{q}]$ and $j\leq J$. Here all other elements of $\theta_{j}^{*}$ are zero and $d\geq J^{q}$ so $\|\bm{\theta}^{*}\|_{0}=J$. The observations are sampled as $y_{i}\sim N(\theta_{i}^{*},\sigma^{2})$. This setting provides a flexible way to control the alignment between the signal structure and the spectrum. When $q=1$, the ordering of $\bm{\theta}^{*}$ align perfectly with the ordering of $\bm{\lambda}$. As $q$ increases, more nonzero elements of $\bm{\theta}^{*}$ are located on the tail where the eigenvalues are smaller, and more large eigenvalues are associated with zero signals, creating a worse signal-spectrum alignment.
@@
The first experiment visualizes the span profile of the signal w.r.t. the learned spectrum at various stages in the OP-GF process with $D=0$. Given a sample, we approximate the gradient flow in Equation˜13 by discrete-time gradient descent and obtain the solution $\{(a_{j}(t),\beta_{j}(t))_{t\geq 0}:j\in[d]\}$. The trained eigenvalue sequence $\tilde{\bm{\lambda}}(t)$ at time $t$ is given by $\tilde{\lambda}_{j}(t)=a_{j}^{2}(t)$ for $j\in[d]$. Here we focus on time points before the optimal stopping time. Figure 1 illustrates the evolution of the span profile w.r.t. the learned spectrum for different training times $t$ and various values of $q$.
@@
We next empirically investigate the evolution of the ESD $d^{\dagger}$ and the estimation error as well as the impact of layers $D$. At any time $t$, we compute the ESD $d^{\dagger}(t)$ based on the learned eigenvalue sequence $\tilde{\bm{\lambda}}(t)$ $\tilde{\bm{\lambda}}(t)$ and also the PC estimate $\widehat{\bm{\theta}}(t)$ based on $\tilde{\bm{\lambda}}(t)$, with number of components determined by $d^{\dagger}(t)$. Theorems˜3.2 and 3.3 suggest that the PC estimator tuned by the ESD can achieve the minimax risk rate, so we expect $\widehat{\bm{\theta}}(t)$ to perform well.
@@
The empirical evaluation involved 20 Monte Carlo repetitions. Figure˜2 displays the averaged $d^{\dagger}(t)$ and the averaged estimation error of $\widehat{\bm{\theta}}(t)$ in Figure˜2 as a function of training time $t$. We observe that both the ESD and the squared error of the tune PC estimator exhibit a general decay trend over training time $t$. Furthermore, for the shallow model with $D=0$ (with no $b_{i,j}$ parameters), the initial decrease in ESD and MSE occurs earlier compared to the deeper models with $D=1$ or $D=3$. However, with sufficient training iterations, the deeper models with $D=1$ or $D=3$ can achieve lower ESD values than the shallow model with $D=0$. These findings suggest that increased model depth ($D>0$) may facilitate a better adaptation of the spectrum, and thus lead to lower estimation error. This observation offers a perspective on the benefits of depth in spectral learning, but a comprehensive study for general models is left for future research.

### v3

@@ 3.2 Span Profile
The span profile $\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}}$ is a well-defined object that depends only on $\bm{\theta}^{*}$ and the ordering of $\bm{\lambda}$, and it summarizes how $\sigma^{2}$ affects the ESD. Theorem˜3.2 suggests that for two spectra $\bm{\lambda}^{(1)}$ and $\bm{\lambda}^{(2)}$, we can compare their alignments with the signal by the ratio of $r(\tau)={\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}^{(1)}}(\tau)}/{\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}^{(2)}}(\tau)}$ for small $\tau$, because, if this ratio is very small (and in particular if the limit is 0 for $\tau\to 0$), then a kernel method using $\bm{\lambda}^{(1)}$ can achieve a smaller risk than one that uses $\bm{\lambda}^{(2)}$. Such comparisons are not as convenient in classical theory. See Section˜10 for more illustrations.
A closely related object is the trade-off function of $\bm{\theta}^{*}$ relative to $\bm{\lambda}$, which is defined for any $k\in[d]$ as
$\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}}(k)\;=\;\frac{1}{k}\sum_{i=k+1}^{d}\left(\theta_{\pi_{i}}^{*}\right)^{2}\;=\;\frac{1}{k}\sum_{\,i:\,\lambda_{i}<\lambda_{\pi_{k}}\,}\left(\theta_{i}^{*}\right)^{2}.$ (9)
@@
Property (3) in Proposition˜3.6 suggests that the faster $\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}}(\cdot)$ decreases, the better the spectrum $\bm{\lambda}$ aligns with the signal $\bm{\theta}^{*}$. In the extreme case where the ordering of $\lambda_{i}$ matches the ordering of $|\theta_{i}^{*}|^{2}$}, the decay of $\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}}(\cdot)$ is the fastest, which leads to the most favorable span profile.
Extensions.
To save space, we defer the extensions to linear models and kernel regression to Sections˜8 and 9, respectively. For the kernel regression model in Equations˜1 and 2, we define the ESD of $f^{*}$ w.r.t. the kernel $\mathbf{k}$ and the effective noise variance $\sigma^{2}:=(\sigma_{0}^{2}+\|f^{*}\|_{\infty}^{2})/n$ as
$d^{\dagger}(\sigma^{2};f^{*},\mathbf{k})=\min\{k\in\mathbb{N}_{+}\cup\{\infty\}:\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}}(k)\leq\sigma^{2}\}.$
@@ 4 Minimax Optimal Convergence Rates
(1) Suppose $K_{n}\asymp n^{a}$ where $0<a<1$. For any $k$, we have $M_{k}\asymp k^{1/a}$. Since $k/k^{1/a}$ is decreasing, ˜4.1 holds.
(2) Suppose $K_{n}\asymp(\log n)^{b}$ where $b>0$. For any $k$, we have $M_{k}\asymp e^{k^{1/b}}$. Since $k/e^{k^{1/b}}$ is decreasing, ˜4.1 holds.
The next theorem provides a lower bound on the minimax risk in Equation˜11.
@@
Connecting adaptivity and generalization. Our span profile framework clarifies why adaptive machine learning methods often outperform classical fixed-kernel approaches. In classical methods with a fixed kernel spectrum $\bm{\lambda}^{(0)}$, the target signal $\bm{\theta}^{*}$ might exhibit poor alignment, resulting in a large span profile $\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}^{(0)}}$. Consequently, the signal resides in a class $\mathcal{F}_{\mathbf{K}^{(0)},\bm{\lambda}^{(0)}}$ with a large quota sequence $\mathbf{K}^{(0)}$, which implies high minimax risk. By contrast, adaptive methods modify the kernel during training. Successful adaptation improves the alignment by reducing the span profile $\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}^{(a)}}$ of the same signal w.r.t. the adapted kernel spectrum $\bm{\lambda}^{(a)}$. This adaptation places the signal in a class $\mathcal{F}_{\mathbf{K}^{(a)},\bm{\lambda}^{(a)}}$ with a smaller quota sequence $\mathbf{K}^{(a)}$, which implies lower minimax risk.
5 Why ridge cannot define an intrinsic complexity measure
This section explains why a complexity measure defined by using the ridge estimator as a guideline is not suitable for characterizing intrinsic generalization difficulty. The key issue is the classical saturation effect of ridge: ridge cannot fully exploit the smoothness property of signals that are sufficiently smooth, so its optimal bias–variance trade-off becomes insensitive to the true source class.
5.1 Ridge estimator and risk decomposition
We work with the same Gaussian sequence model as in Equation˜4. The ridge estimator can be written in the generic spectral form $\widehat{\theta}^{\mathrm{R},\nu}_{j}=\bigl(1-\psi_{\nu}^{\mathrm{R}}(\lambda_{j})\bigr)z_{j}$ with $\psi_{\nu}^{\mathrm{R}}(\lambda)=\nu/(\lambda+\nu)$, namely
$\widehat{\theta}^{\mathrm{R},\nu}_{j}=\frac{\lambda_{j}}{\lambda_{j}+\nu}\,z_{j}.$
Its mean squared error admits the standard bias–variance decomposition
$\mathcal{R}_{\mathrm{R}}(\nu)=\mathbb{E}\Bigl[\|\widehat{\bm{\theta}}^{\mathrm{R},\nu}-\bm{\theta}^{*}\|_{2}^{2}\Bigr]=B_{\mathrm{R}}(\nu)+V_{\mathrm{R}}(\nu),$
where the squared bias is given by
$B_{\mathrm{R}}(\nu)=\sum_{j=1}^{d}\Bigl(\psi_{\nu}^{\mathrm{R}}(\lambda_{j})\Bigr)^{2}(\theta^{*}_{j})^{2}=\sum_{j=1}^{d}\Bigl(\frac{\nu}{\lambda_{j}+\nu}\Bigr)^{2}(\theta^{*}_{j})^{2},$
and the variance is
$V_{\mathrm{R}}(\nu)=\sigma^{2}\sum_{j=1}^{d}\Bigl(1-\psi_{\nu}^{\mathrm{R}}(\lambda_{j})\Bigr)^{2}=\sigma^{2}\sum_{j=1}^{d}\Bigl(\frac{\lambda_{j}}{\lambda_{j}+\nu}\Bigr)^{2}.$
Before moving on, we introduce the ordering and a convenient reparametrization. Given the spectrum $\bm{\lambda}=\{\lambda_{j}\}_{j=1}^{d}$, let $\pi$ be the permutation that sorts the eigenvalues in decreasing order: $\lambda_{\pi_{1}}>\lambda_{\pi_{2}}>\cdots$ (assume no ties for simplicity), and write $\lambda_{(i)}:=\lambda_{\pi_{i}}$. For any $\nu\in(0,\lambda_{(1)}]$, define the index
$k_{\bm{\lambda}}(\nu)=\#\{j:\lambda_{j}\geq\nu\}.$
In words, $k_{\bm{\lambda}}(\nu)$ is the number of eigenvalues no smaller than $\nu$. By definition, we have $\lambda_{(k_{\bm{\lambda}}(\nu))}\geq\nu>\lambda_{(k_{\bm{\lambda}}(\nu)+1)}.$
5.2 Lower bound on the variance
We introduce a convenient lower bound for the variance term $V_{\mathrm{R}}(\nu)$.
For any $k\in[d]$, define
$\widetilde{N}(k,d;\bm{\lambda})=\frac{\sum_{i=k+1}^{d}\lambda_{(i)}^{2}}{k\,\lambda_{(k)}^{2}},$
and define
$\widetilde{V}(k):=\sigma^{2}\,k\bigl(1+\widetilde{N}(k,d;\bm{\lambda})\bigr)=\sigma^{2}\,\left(k+\frac{\sum_{i=k+1}^{d}\lambda_{(i)}^{2}}{\,\lambda_{(k)}^{2}}\right).$
Proposition 5.1.
Fix $\nu\in(0,\lambda_{(1)}]$ and let $k:=k_{\bm{\lambda}}(\nu)$. It holds that
$\frac{1}{4}\widetilde{V}(k)\leq V_{\mathrm{R}}(\nu).$
5.3 A lower bound on the ridge bias at small regularization
The saturation phenomenon is driven by the fact that for small $\nu$ the ridge bias behaves as
$\nu^{2}\sum_{j}(\theta^{*}_{j}/\lambda_{j})^{2},$
which involves the rescaled coefficients $\theta^{*}_{j}/\lambda_{j}$ and therefore can stop decreasing once the signal is smoother than ridge can exploit.
To be concrete, we place the following assumption.
Assumption 5.2.
There exist an index $i_{0}\in[d]$ and a constant $c_{r}>0$ such that
$\sum_{i=1}^{i_{0}}\frac{(\theta^{*}_{\pi_{i}})^{2}}{\lambda_{(i)}^{2}}\geq 4c_{r}.$
We have the following lower bound on the squared bias term $B_{\mathrm{R}}(\nu)$
Proposition 5.3.
Under ˜5.2, it holds for any $\nu\leq\lambda_{(i_{0})}$ that
$B_{\mathrm{R}}(\nu)\geq c_{r}\,\nu^{2}.$ (12)
For any $k$, we define the proxy for the bias as
$\widetilde{B}(k):=c_{r}\,\lambda_{(k)}^{2}.$
Note that $\widetilde{B}(k)$ is comparable to the lower bound $c_{r}\nu^{2}$ if $k=k_{\bm{\lambda}}(\nu)$ and $k\geq i_{0}$.
5.4 A ridge-guided complexity.
Suppose ˜5.2 holds. Combining the bias proxy $\widetilde{B}(k)$ and the variance proxy $\widetilde{V}(k)$, we consider the proxy risk
$\widetilde{R}(k)=\widetilde{B}(k)+\widetilde{V}(k)=c_{r}\,\lambda_{(k)}^{2}+\sigma^{2}\,k\bigl(1+\widetilde{N}(k,d;\bm{\lambda})\bigr).$
This leads to the ridge-guided analogue of the trade-off sequence used in the definition of ESD. Specifically, we define the ridge trade-off sequence
$\bar{h}(k;\bm{\theta}^{*},\bm{\lambda})=\frac{\lambda_{(k)}^{2}}{k\bigl(1+\widetilde{N}(k,d;\bm{\lambda})\bigr)}.$
Definition 5.4.
Define the ridge saturating dimension by
$d^{\Delta}(\bm{\theta}^{*},\bm{\lambda},\sigma^{2})=\max\Bigl\{k:\ c_{r}\,\bar{h}(k;\bm{\theta}^{*},\bm{\lambda})>\sigma^{2}\Bigr\}.$
By construction, at $k=d^{\Delta}$ the two proxy terms are balanced up to constants:
$c_{r}\,\lambda_{(d^{\Delta})}^{2}>\sigma^{2}\,d^{\Delta}\bigl(1+\widetilde{N}(d^{\Delta},d;\bm{\lambda})\bigr),\qquad c_{r}\,\lambda_{(d^{\Delta}+1)}^{2}\leq\sigma^{2}\,(d^{\Delta}+1)\bigl(1+\widetilde{N}(d^{\Delta}+1,d;\bm{\lambda})\bigr).$
It turns out, we can use $d^{\Delta}$ to lower bound the risk of the ridge estimator. The key point is that $d^{\Delta}$ is the quantity that ridge cannot beat (up to constants) when $\nu$ ranges over the natural regime $\nu\leq\lambda_{(i_{0})}$.
Proposition 5.5.
Suppose ˜5.2 holds. For any integer $k\geq i_{0}$, it holds that
$\widetilde{R}(k)\geq\sigma^{2}\,d^{\Delta}(\bm{\theta}^{*},\bm{\lambda},\sigma^{2}).$
The proof is straightforward:
  • if $k\geq d^{\Delta}$, then $\widetilde{R}(k)\geq\widetilde{V}(k)\geq\sigma^{2}k\geq\sigma^{2}d^{\Delta}$;
  • if $k\leq d^{\Delta}$, then $\widetilde{R}(k)\geq\widetilde{B}(k)\geq\widetilde{B}(d^{\Delta})>\sigma^{2}d^{\Delta}$.
Theorem 5.6 (Lower bound for Ridge risk).
Suppose ˜5.2 holds. For any $\nu\in(0,\lambda_{(i_{0})}]$, let $k:=k_{\bm{\lambda}}(\nu)$. Then
$\mathcal{R}_{\mathrm{R}}(\nu)\geq\max\Bigl\{c_{r}\nu^{2},\ \frac{1}{4}\widetilde{V}(k)\Bigr\}\geq\frac{1}{4}\sigma^{2}\,d^{\Delta}(\bm{\theta}^{*},\bm{\lambda},\sigma^{2}).$
Consequently,
$\inf_{\nu\in(0,\lambda_{(i_{0})}]}~\mathcal{R}_{\mathrm{R}}(\nu)\geq\frac{1}{4}\sigma^{2}\,d^{\Delta}(\bm{\theta}^{*},\bm{\lambda},\sigma^{2}).$
5.5 Why this ridge-guided index fails to recover source-class structure
As discussed in the main text (see Theorem˜3.3), the minimax risk over the class
$\mathcal{F}_{K,\bm{\lambda}}^{(n)}=\Bigl\{\bm{\theta}\in\mathbb{R}^{d}:d^{\dagger}(\sigma^{2};\bm{\theta},\bm{\lambda})\leq K\Bigr\}$ (13)
scales as $\sigma^{2}K$ up to constants. Therefore, $d^{\dagger}$ captures the intrinsic difficulty of estimation.
By contrast, the risk of ridge estimator is lower bounded by $d^{\Delta}$, which depends on $\bm{\lambda}$ through $\lambda_{(k)}^{2}$ and $\widetilde{N}(k,d;\bm{\lambda})$, and it becomes insensitive to additional smoothness once the signal is beyond ridge’s qualification. In particular, we can use the ratio
$\frac{d^{\dagger}(\bm{\theta}^{*},\bm{\lambda},\sigma^{2})}{d^{\Delta}(\bm{\theta}^{*},\bm{\lambda},\sigma^{2})}$
to study whether the saturation effect appears for any given signal and spectral. In particular, if this ratio is very small, Theorem˜9.6 and Theorem˜5.6 together suggests the risk of ridge estimator is much larger than the risk of PC estimator. We demonstrate this idea using the following example.
Example 5.7.
Suppose ˜5.2 holds and consider the classical polynomial spectrum that $\lambda_{(i)}=i^{-\beta}$ where the degree $\beta>1/2$. It can be showed that $\widetilde{N}(k,d;\bm{\lambda})$ is bounded by a constant for all $k$. Therefore, we have
$\bar{h}(k;\bm{\theta}^{*},\bm{\lambda})\asymp\frac{\lambda_{(k)}^{2}}{k}=\frac{k^{-2\beta}}{k}=k^{-(1+2\beta)}.$
Given any $\sigma^{2}>0$, we have
$d^{\Delta}(\bm{\theta}^{*},\bm{\lambda},\sigma^{2})\gtrsim(\sigma^{2})^{-1/(1+2\beta)}.$
Now consider a classical source condition: there exists $R>0$ and $s>0$ such that
$\sum_{i=1}^{d}\lambda_{i}^{-s}(\theta_{i}^{*})^{2}\leq R.$
This implies
$h(k;\bm{\theta}^{*},\bm{\lambda})=\frac{1}{k}\sum_{i=k+1}^{d}(\theta^{*}_{\pi_{i}})^{2}\leq\frac{R}{k}\,\lambda_{(k)}^{s}=R\,k^{-(1+s\beta)},$
and hence
$d^{\dagger}(\bm{\theta}^{*},\bm{\lambda},\sigma^{2})\lesssim(\sigma^{2})^{-1/(1+s\beta)}.$
If $s>2$ (i.e., the signal is smoother than ridge’s qualification), then
$\frac{d^{\dagger}(\bm{\theta}^{*},\bm{\lambda},\sigma^{2})}{d^{\Delta}(\bm{\theta}^{*},\bm{\lambda},\sigma^{2})}\ \lesssim\ (\sigma^{2})^{\frac{1}{1+2\beta}-\frac{1}{1+s\beta}}\ \to 0\qquad\text{as }\sigma^{2}\to 0,$
so the ridge-guided index $d^{\Delta}$ is asymptotically much larger than the intrinsic complexity $d^{\dagger}$. Equivalently, ridge is forced to pay a risk on the order of $\sigma^{2}d^{\Delta}$, even though the minimax rate over the same source class continues to improve with $s$ and is on the order of $\sigma^{2}d^{\dagger}$. ∎
The theory developed here and the above example demonstrate that when the signal is sufficiently smooth, ridge cannot translate that additional smoothness into a smaller optimal bias term, and the resulting ridge-guided index complexity index $d^{\Delta}$ becomes defective to characterize the optimal risk. This is why it is unsuitable as a replacement for ESD in our rate statements.
@@ 6 Adaptive Eigenvalues via Over-parameterized Gradient Flow
This section will investigate the benefits of learning eigenvalues via over-parameterized gradient flow (OP-GF) in sequence models [29] through the lens of ESDs. For more general adaptive kernels with evolving eigenfunctions, Section˜10.2 outlines how ESD can be used as an explanatory framework and we provide supporting evidence via a linear network experiment. Analyzing eigenfunction evolution technically more difficult than eigenvalues evolution and is left for future work.
@@
Suppose that ˜6.1 holds and the initialization in Equation˜14 is $b_{0}=c_{B}D^{\frac{D+1}{D+2}}\varepsilon^{\frac{1}{D+2}}$. Define $t_{2}=C\cdot D^{\frac{D}{D+2}}(\varepsilon)^{-\frac{2D+2}{D+2}}$. There exist some constants $c$, $C$, $C_{M}$, $C_{max}$, $C_{\eta}$, $c_{\eta}$, $c_{B}$, and $c^{\prime}$, such that with probability larger than 14/n1-4/n, we have
@@
Theorem˜6.2 shows that OP-GF reduces the ESD. The conditions ensure that strong signals with small initial eigenvalues have larger adapted eigenvalues than weak signals.
@@ 7 Numerical experiments
Data Generation. We utilize the misalignment setting in Li and Lin [29] to specify a $d$-dimensional sequence model. We fixed the eigen-decay rate $\gamma>0$, the signal decay rate $p>0$, and the number of nonzero signals $J$. Given any misalignment parameter $q\geq 1$, we set eigenvalues as $\lambda_{j}=j^{-\gamma},j\in[d]$, and set the true nonzero parameters as $\theta_{\ell(j)}^{*}=C\cdot j^{-\frac{p+1}{2}}$, where $\ell(j)=[j^{q}]$ and $j\leq J$. Here all other elements of $\theta_{j}^{*}$ are zero and $d\geq J^{q}$ so $\|\bm{\theta}^{*}\|_{0}=J$. The observations are sampled as $y_{i}\sim N(\theta_{i}^{*},\sigma^{2})$. This setting provides a flexible way to control the alignment between the signal structure and the spectrum. When $q=1$, the ordering of $\bm{\theta}^{*}$ aligns perfectly with the ordering of $\bm{\lambda}$. As $q$ increases, more nonzero elements of $\bm{\theta}^{*}$ are located on the tail where the eigenvalues are smaller, and more large eigenvalues are associated with zero signals, creating a worse signal-spectrum alignment.
@@
The first experiment visualizes the span profile of the signal w.r.t. the learned spectrum at various stages in the OP-GF process with $D=0$. Given a sample, we approximate the gradient flow in Equation˜14 by discrete-time gradient descent and obtain the solution $\{(a_{j}(t),\beta_{j}(t))_{t\geq 0}:j\in[d]\}$. The trained eigenvalue sequence $\tilde{\bm{\lambda}}(t)$ at time $t$ is given by $\tilde{\lambda}_{j}(t)=a_{j}^{2}(t)$ for $j\in[d]$. Here we focus on time points before the optimal stopping time. Figure 1 illustrates the evolution of the span profile w.r.t. the learned spectrum for different training times $t$ and various values of $q$.
@@
We next empirically investigate the evolution of the ESD $d^{\dagger}$ and the estimation error as well as the impact of layers $D$. At any time $t$, we compute the ESD $d^{\dagger}(t)$ based on the learned eigenvalue sequence $\tilde{\bm{\lambda}}(t)$ and also the PC estimate $\widehat{\bm{\theta}}(t)$ based on $\tilde{\bm{\lambda}}(t)$, with number of components determined by $d^{\dagger}(t)$. Theorems˜3.2 and 3.3 suggest that the PC estimator tuned by the ESD can achieve the minimax risk rate, so we expect $\widehat{\bm{\theta}}(t)$ to perform well.
@@
The empirical evaluation involved 20 Monte Carlo repetitions. Figure˜2 displays the averaged $d^{\dagger}(t)$ and the averaged estimation error of $\widehat{\bm{\theta}}(t)$ in Figure˜2 as a function of training time $t$. We observe that both the ESD and the squared error of the tuned PC estimator exhibit a general decay trend over training time $t$. Furthermore, for the shallow model with $D=0$ (with no $b_{i,j}$ parameters), the initial decrease in ESD and MSE occurs earlier compared to the deeper models with $D=1$ or $D=3$. However, with sufficient training iterations, the deeper models with $D=1$ or $D=3$ can achieve lower ESD values than the shallow model with $D=0$. These findings suggest that increased model depth ($D>0$) may facilitate a better adaptation of the spectrum, and thus lead to lower estimation error. This observation offers a perspective on the benefits of depth in spectral learning, but a comprehensive study for general models is left for future research.